# A Closed-Form Solution for Kernel Adaptive Filtering

### v2

@@ Abstract
Conventional kernel adaptive filtering (KAF) uses a prescribed, positive-definite, nonlinear function to define the Reproducing Kernel Hilbert Space (RKHS), where the minimum mean square error solution is found using search techniques. This paper proposes embedding the statistics of the input data in the kernel definition, obtaining a closed-form solution for nonlinear adaptive filtering applications. We call this solution the Functional Wiener Filter (FWF), which extends Parzen’s work on the autocorrelation RKHS to nonlinear functional spaces. We first derive a space-time covariance operator on a RKHS that generalizes Parzen’s autocovariance RKHS and includes nonlinear functions of a random process. We then present a method for approximating the FWF in an explicit, finite-dimensional RKHS to model time series directly from realizations. The FWF can be computed in closed form directly from realizations and has better accuracy than other kernel-based algorithms in a synthetic strictly stationary dataset. We show that FWF performance is comparable to other KAF algorithms for two chaotic time series and a real-world time series. The FWF’s computational complexity at test time is an improvement over other KAFs. We demonstrate how the difference equation learned by the FWF can be extracted, leading to system identification applications, beyond optimal nonlinear filtering.
@@ 1 Introduction
In 1949, Norbert Wiener published his book, ”Extrapolation, Interpolation, and Smoothing of Stationary Time Series” [38], where he detailed his seminal work on linear minimum mean squared error (MMSE) estimation. While this work launched statistical modeling of time series, the Wiener-Hopf method [39] is rather complex and the optimal solution still belongs to the span of the input data i.e., the corresponding model is linear.
Briefly, let $X$ be a wide-sense multivariate stationary random process, and $Z$ be a target measurement of a multivariate random variable. Define the error $e=Z-\hat{Z}$, where $\hat{Z}$ is any arbitrary function of $X$ and $Z$, called the estimator. By definition, the mean square error (MSE) is the trace of the autocorrelation of the error $\mathbb{E}[ee^{\top}]$. The MMSE estimator finds the $\hat{Z}$ which minimizes this trace. Alternatively, the minimal MSE can be found using gradient descent techniques, with the advantage of allowing nonlinear mapping functions of $X$, but the solution is neither closed-form nor unique. In a Bayesian setting, the MMSE estimator is given by the posterior mean. Since the posterior mean involves an integration, the form of the MMSE estimator is usually constrained to be within a certain class of functions if a closed-form solutions is desired. Linear MMSE estimators are in widespread use in statistics and signal processing in part because they allow for a closed-form solution and provide a simple interpretation of the learned model. The main contribution of this paper is the construction of a reproducing kernel Hilbert space (RKHS) to solve the posterior mean analytically as a minimum norm projection, yielding a closed-form solution for nonlinear MMSE estimation that differs from the ones available in the literature.
In the late 50s, Emmanuel Parzen [25] presented an alternative approach to solve the linear MMSE problem in a RKHS defined by the autocovariance function of a wide-sense stationary random process (r.p.). He shows that the MMSE solution in this RKHS is exactly the inner product of the input time series with the cross-covariance function between the input time series and the desired time series.
Parzen argues that the embedding of the statistical information of the autocovariance function into the inner product of the RKHS creates a natural space for statistical inference on time series because conditional expectations with respect to the stationary input r.p. can be expressed as inner products in the space (under some Gaussian assumptions). This claim of a “natural” space is supported by the subsequent use of the autocovariance RKHS($\mathcal{H}_{R}$) to clarify and simplify many problems such as the detection of deterministic signals in colored Gaussian noise [13], detection of Gaussian signals in Gaussian noise [12], generalized innovations representations [11], calculation of the likelihood-ratio in Gaussian [11] and non-Gaussian settings [4], and parameter estimation [5].
This paper takes Parzen’s theoretical work one step further by deriving a closed-form solution for a nonlinear MMSE estimator in a data-dependent RKHS. Inspired by [25], we also project the input data to a RKHS, but now the kernel is data-dependent and nonlinearly related to the input space. While Kernel Adaptive Filtering (KAF) [17] poses the problem of nonlinear MMSE filter design as a search in an RKHS defined by a data-independent, nonlinear kernel, the proposed method is quite different. It first designs a data-dependent universal RKHS and then computes, in closed-form, the minimum norm projection of the target in the constructed space, avoiding gradient search. While a closed-form solution to select parameters of a nonlinear filter for regression is foundational in its own right, the method for embedding the statistics of input data, i.e. its internal dependency structure, into the inner product of an RKHS defined by a nonlinear kernel is perhaps the more generally applicable contribution of this work.
To achieve our goal, we show that full access to individual eigenfunctions of the kernel is necessary, which is beyond the reach of the conventional “kernel trick” [31]. Instead, for the implementations, we approximate the Gaussian kernel by an explicit, finite-dimensional RKHS where the inner product becomes an explicit quadratic form of vectors. The disadvantage is that the explicit RKHS is finite-dimensional i.e., no longer universal or characteristic. However, the previous results in [14] show that only a few dimensions are needed to achieve excellent performance with KAF algorithms.
The main contributions of this work are the following:
  • We introduce an extension of Parzen’s MMSE in $\mathcal{H}_{R}$ to an RKHS that includes nonlinear functions of a time series, yielding a closed-form, computationally efficient, and effective nonlinear MMSE estimator.
  • We provide a method for the practical implementation and use of a data-dependent nonlinear RKHS for signal processing applications.
  • We show experimentally that if the assumptions made in the closed-form solution are met we outperform other kernel-based nonlinear filtering methods.
  • We show that the optimal solution parameters can be interpreted as a nonlinear difference equation learned directly from the data. This extends possible applications of the method to system identification tasks and physics based modeling in RKHS.
We will highlight some more related works next. In [37] a comprehensive treatment of spline models is given. The problem of finding an optimal MMSE spline to fit data is presented in an RKHS, however, the formulation of the solution is quite different from ours in that our solution first focuses on defining a nonlinear data-dependent space where the MMSE solution can be immediately given. In [4] the idea of a covariance function is extended to a characteristic functional, which under certain conditions can characterize non-Gaussian r.p.s. (see [9] for a description of the characteristic functional). This is related to our work in that it seeks to define a data-dependent RKHS. While the results in [4] are theoretically interesting, the conditions are restrictive, and its application to general problems such as MMSE filtering is not clear.
@@
More contemporary related work is centered on Kernel Adaptive Filtering (KAF). A comprehensive description of the foundational work can be found in [17]. In general, KAF utilizes gradient search methods to find the optimal functional in the feature space of an RKHS resulting in a nonlinear filter. Some examples of this approach are Kernel Least Mean Squares (KLMS) [16] and Kernel Recursive Least Squares [6] which extend the least mean squares and recursive least squares algorithms respectively.
The strength of many KAF methods is that they are convex universal learning machines with simple implementations. One of the main issues with KAF is that the model size grows with the number of training samples. This problem is addressed by sparsification algorithms such as Quantized Kernel Least Mean Squared (QKLMS) [1]. A comparative study of different KAF techniques with sparsification can be found in [36]. Another notable strategy for decoupling model size with the number of training samples is to use finite-dimensional approximations of the feature space specified by some reproducing kernel. Examples of this can be found in [14], [28], and [18]. In our work, we utilize this finite-dimensional approximation, which decouples model size from training set size. Our method is distinct from other KAF methods in a few consequential ways. Firstly, we eliminate the need for an iterative gradient-based search through an RKHS, hence our claim of a closed-form solution for KAF. Second, while other KAF methods define a functional based solely on the amplitude of a time series, the FWF yields functionals over both time and amplitude values. Third, the FWF focus on first building a data-dependent RKHS where the MMSE solution can immediately be given. This is a fundamentally different approach to solving for a MMSE solution.
@@
Another popular kernel-based regression method is Gaussian Process Regression (GPR) [29]. This method uses a Gaussian Process (a time series where all joint distributions are Gaussian) to define a distribution over possible functions. Then the data is used to update this prior distribution using maximum a posteriori (MAP) estimators. GPR offers a Bayesian framework for nonlinear regression, which is distinct from the adaptive filtering-based framework of KAF.
@@
In [23], the kernel autocovariance operators of stationary processes are theoretically studied, and classical limit theorems as well as non-asymptotic error bounds under some ergodic assumptions are discussed. The authors also give a conditional mean operator based on the autocovariance operator. The autocovariance operators discussed in [23] are closely related to the $U$ operator discussed in later sections of this text. However, [23] is focused on the analysis of the kernel autocovariance operator itself, while our work was independently developed and progresses in the opposite direction. It starts with the foundational ideas of statistical signal processing developed by Wiener, Parzen, and KAF and extends all these concepts to potentially universal data-dependent RKHS and covariance operators. We believe that our approach will be easier to understand by experts in signal processing. Moreover, it provides algorithms for practical applications that can be easily understood and employed by practitioners.
@@ 2 MMSE Solutions in Data-Dependent RKHSs
2.1 Random Processes
Let $(\Omega,\mathcal{A},P)$ be a probability space with sample space $\Omega$, $\sigma$-algebra $\mathcal{A}$, and probability measure $P$. A random process (r.p.), $X=\{X_{t},t\in T\}$, is a collection of random variables (r.v.s.) defined on, $(\Omega,\mathcal{A},P)$, along with an index set $T$ that is a compact subset of a separable metric space usually representing time. All r.p.s. will be assumed to contain real-valued r.v.s.. We denote a single r.v. within a r.p. with a capital letter subscripted with its index, $X_{t}$. Realizations from these random variables will be denoted as $x_{t}$.
In our treatment of the linear MMSE solution, we assume wide-sense stationarity of the processes involved. A wide-sense stationary random process satisfies the following relation [33],
$R(\tau)=\mathbb{E}[X_{t}X_{t-\tau}]$ (1)
for any $t\in T$ and $\tau$ such that $t-\tau\in T$. Likewise, for two jointly stationary random processes $Z$ and $X$,
$\rho(\tau)=\mathbb{E}[Z_{t}X_{t-\tau}]$ (2)
for any $t\in T$ and $\tau$ such that $t-\tau\in T$. In other words, the joint second-order statistics are independent of $t$. This assumption of wide-sense stationarity is necessary to practically estimate the necessary quantities needed to solve the Wiener-Hopf equations from realizations of a random process [38]. Another important notion of stationarity is strict stationarity. A strictly stationary random process is a r.p. where the joint distributions(not just the second-order moments) do not change with shifts in time. This strict stationarity assumption will become important in the nonlinear MMSE solution.
@@ 2.2 Linear MMSE Solution
We now review the linear MMSE solution given in [25]. A more detailed accounting can be found in [26]. Let the space of square-integrable r.v.s defined on $(\Omega,\mathcal{B},P)$ be denoted as $L^{2}(\Omega,\mathcal{A},P)$. This is the space of all r.v.s, $W$, such that,
$\|W\|_{2}=\int_{\Omega}|W|^{2}dP<\infty$ (3)
The linear span of a r.p., $X$, in $L^{2}(\Omega,\mathcal{A},P)$ is the smallest subset of $L^{2}(\Omega,\mathcal{A},P)$ containing $X$ which possess the properties of a Hilbert space [26] (a subspace). We can define this set by first defining the linear manifold $L(X_{t},t\in T)$ as the set of all r.v.s. with the following form
$W=\sum_{i=1}^{n}a_{i}X_{t_{i}}\quad a_{i}\in\mathbb{R},\quad t_{i}\in T,\quad n\in\mathbb{N}$ (4)
While $L(X_{t},t\in T)$ is a linear manifold, it is not complete. We can complete this space by including the limits of all Cauchy sequences of elements in $L(X_{t},t\in T)$. This complete set is the linear span of a r.p. in $L^{2}(\Omega,\mathcal{A},P)$, denoted as $L^{2}(X)$. Consider two r.v.s. $W,V\in L^{2}(X)$ where $W=\sum_{t\in T}a_{t}X_{t}$, $V=\sum_{s\in T}b_{s}X_{s}$, and $a_{t},b_{s}\in\mathbb{R}$. The inner product in $L^{2}(X)$ can be written as,
$\langle W,V\rangle_{L^{2}}=\mathbb{E}[WV]=\mathbb{E}\left[\sum_{s,t\in T}a_{t}b_{s}X_{t}X_{s}\right]=\sum_{s,t\in T}a_{t}b_{s}\mathbb{E}[X_{t}X_{s}]$ (5)
We can then define the covariance function of the r.p. as
$R(s,t)=\mathbb{E}[X_{t}X_{s}]$ (6)
[…]
$\mathbb{E}[WV]=\sum_{s,t\in T}a_{t}b_{s}R(s,t)$ (7)
Since the covariance function is a positive semi-definite function, it defines a reproducing kernel, and thus, we can conclude that (7) is an inner product between two elements in the RKHS defined by the autocovariance function ($\mathcal{H}_{R}$).
$\langle W,V\rangle_{L^{2}}=\sum_{s,t\in T}a_{t}b_{s}R(s,t)=\langle W^{{}^{\prime}},V^{{}^{\prime}}\rangle_{\mathcal{H}_{R}}$ (8)
@@
Suppose we are given the r.p. $X$ as input and the r.p. $Z$ as the desired r.v.. As a consequence of Hilbert projection theorem, the linear MMSE solution can be given as the projection of $Z$ into $H_{R}$ which is simply the crosscovariance function. This projection is given as,
$\rho(s)=\mathbb{E}[ZX_{s}]$ (9)
The MMSE solution is the inner product of $\rho$ with $X$ in $\mathcal{H}_{R}$ that is
$Z^{*}=\langle\rho,X\rangle_{\mathcal{H}_{R}}=\sum_{s,t\in T}X_{t}R(s,t)^{\dagger}\rho(s)$ (10)
Where $\dagger$ is the Moore-Penrose pseudo inverse. In discrete time, this solution is equivalent to the Wiener solution, where the optimal impulse response is given by $w^{*}=\mathbf{R}^{\dagger}\bm{\rho}$, where $\mathbf{R}$ the auto-covariance matrix and $\bm{\rho}$ is the crosscovariance vector. In [25], it is shown in greater detail that the cross-covariance function is the linear MMSE solution in $\mathcal{H}_{R}$. This demonstration shows that with the data-dependent RKHS, $\mathcal{H}_{R}$, there is no need to search for the linear MMSE solution. The solution is the cross-covariance function. The problem is that this solution is still linear in the input space. The natural question that arises is, “can we generalize this idea such that our search space includes nonlinear functions of the input?”. The next section shows the generalization of Parzen’s work in [25].

### v3

@@ Abstract
Unlike the conventional kernel adaptive filtering (KAF) approach of using a fixed kernel to define the Reproducing Kernel Hilbert Space (RKHS), this paper embeds the statistics of the input data in the kernel definition, obtaining a closed-form solution for nonlinear adaptive filtering. We call this solution the Functional Wiener Filter (FWF), and it is formally an extension of Parzen’s work on the autocorrelation RKHS to nonlinear functional spaces. We present a method for approximating the FWF in an explicit, finite-dimensional RKHS to model time series directly from realizations, which is less computationally demanding at test time than other KAF methods. We show that FWF outperforms KAF on a synthetic dataset that meets the conditions of the theory, and is comparable to other KAF algorithms for both a chaotic and real-world time series. We demonstrate how the difference equation learned by the FWF can be extracted, leading to possible applications in system identification.
@@ 1 Introduction
While linear adaptive filters [1] are well-established for applications including system identification, radar and sonar, active noise cancellation, channel equalization, and others, the inherent nonlinearity of most real-world systems motivates the development of nonlinear variants of these algorithms. Despite their flexibility for modeling, in comparison with linear adaptive filtering algorithms, the theory for developing tractable and computationally efficient algorithms for nonlinear adaptive filtering is far from mature. In this work we bridge this gap, by extending the theory for minimum mean square error (MMSE) estimation developed by Emmanuel Parzen, to derive a closed-form solution for nonlinear adaptive filtering.
@@
Kernel adaptive filtering (KAF) [2] offers a solution for nonlinear MMSE filtering. In general, KAF utilizes gradient search methods to find the optimal functional in the feature space of an RKHS, resulting in a nonlinear filter. Some examples of this approach are Kernel Least Mean Squares (KLMS) [3] and Kernel Recursive Least Squares [4], which extend the Least Mean Squares (LMS) and Recursive Least Squares (RLS) algorithms, respectively. KAF is an active area of research, for a comparative study of some KAF methods, see [5]. More recent work on robust KAF includes [6, 7]. Methods for computationally efficient KAF can be found in [8, 9].
While effective and computationally feasible closed-form solutions for KAF have largely been absent, a few prior attempts have been made. For instance, in [10] a nonlinear extension for the Wiener filter based on correntropy [11] was discussed but had non-competitive performance and large test-time computational costs.
@@
Other alternative methods for nonlinear MMSE filtering can be obtained in a Bayesian setting, in which the MMSE estimator is given by the posterior mean. A well-known Bayesian method for nonlinear MMSE estimation is Gaussian Process Regression (GPR) [12]. This method uses a Gaussian process (a random process where all joint distributions are Gaussian) to define a prior distribution over possible functions. Then the data is used to create maximum a posteriori estimators. Other related methodologies are based in the theory of splines [13], where the problem of finding an optimal MMSE spline to fit data is presented in an RKHS.
There are also deep learning based methods, such as LSTMs [14, 15], Temporal Convolutional Neural Networks [16], and Transformers [17]. A recent survey of these methods is given in [18]. However, the computational complexity of these models is not comparable with KAF, and the advances, theoretical and otherwise, made in this paper are most closely related to KAF and GPR. Therefore, we focus on comparison with KAF and GPR.
The theory of optimal linear filtering based on MMSE estimation was initially introduced in the seminal works of Norbert Wiener [19] and Andrey Kolmogorov [20], but a closed-form solution for nonlinear filters has been elusive. In [21], Parzen realized that because the autocorrelation function of the input random process is positive definite it can be used to define a data-dependent reproducing kernel Hilbert space (RKHS), where the optimal filter, in the MMSE sense, corresponds to a linear functional in the space. Parzen argued that embedding the statistical information of the autocovariance function into the inner product of the RKHS creates a natural space for statistical inference on random processes because conditional expectations with respect to a stationary input random process can be expressed as inner products in the RKHS. This claim of a “natural” space is supported by the subsequent use of the autocovariance RKHS, denoted here as $\mathcal{H}_{R}$, to clarify and simplify many problems in statistical signal processing (see [22, 23, 24] for examples).
@@
In [25], the kernel autocovariance operators of stationary processes are theoretically studied and classical limit theorems as well as non-asymptotic error bounds under some ergodic assumptions are discussed. The autocovariance operators discussed in [25] are closely related to the $U$ operator discussed in later sections of this paper. However, [25] is focused on the analysis of the kernel autocovariance operator itself, while our work was independently developed and focuses on the extension of Parzen’s work on MMSE filters.
Our method is closely related to KAF in the way we construct a space of nonlinear functions of the input space, but it connects this space with Parzen’s autocorrelation RKHS to obtain a closed-form solution. Although this closed-form solution requires an assumption that the data are stationary, it remains desirable for several important reasons. First, it gives insight into the system’s performance allowing for more explicit characterization of design variable’s effects on performance metrics. Second, it connects theory to practice, facilitating clearer principles for system design. Third, the optimal solution comes with guarantees; for example, we know that given our simplifying assumptions and our data, the solution is optimal, which aids in the analysis of the underlying system that creates the data. Finally, it may aid the development of optimal control laws. Therefore, the pursuit of a closed-form solution for nonlinear MMSE filters is worthwhile.
Beyond being a closed-form solution, the FWF differs from other KAF methods in several consequential ways. First, while other KAF methods define a functional based solely on the amplitude of a random process, the FWF yields functionals over both time and amplitude values. Second, the FWF focuses on first building a data-dependent RKHS where the MMSE estimator can be given immediately. This fundamentally different approach to solving for the MMSE estimator makes the FWF distinct from other KAF methods.
In summary, the main contributions of this work are the following. First, we introduce the theory to extend Parzen’s MMSE solution in $\mathcal{H}_{R}$ to an RKHS that includes nonlinear functions of a random process, along with a method for computing this solution from realizations, yielding a closed-form, computationally efficient, and effective nonlinear MMSE estimator. Second, we provide a method for the practical implementation and use of a data-dependent nonlinear RKHS for signal processing applications. Third, we demonstrate experimentally that when the assumptions made in the closed-form solution are met, we outperform other kernel-based nonlinear filtering methods. Finally, we show that the optimal solution parameters can be interpreted as a nonlinear difference equation learned directly from the data. This extends possible applications of the method to system identification tasks and physics-based modeling in an RKHS.
The remainder of the paper is organized as follows. First, a review of Parzen’s linear MMSE solution in $\mathcal{H}_{R}$ is given. Next, we introduce the theory that extends this solution to include nonlinear functions. Then, we provide a practical method for computing this solution from realizations of input and target random processes. In sections 3.3-3.4, we give an analysis of the error given by the FWF, and show that the solution given by the FWF can be interpreted as a nonlinear difference equation. Finally, we compare the FWF with other KAF and nonlinear regression methods on two simulated time series and one real-world time series, and give concluding remarks.
@@ 2.1 Linear MMSE Solution
Let $(\Omega,\mathcal{A},P)$ be a probability space with sample space $\Omega$, $\sigma$-algebra $\mathcal{A}$, and probability measure $P$. A random process, $X=\{X_{t},t\in T\}$, is a collection of random variables (r.v.) defined on, $(\Omega,\mathcal{A},P)$, along with an index set $T$ that is a compact subset of a separable metric space usually representing time. All random processes will be assumed to contain real-valued r.v.s. We denote a single r.v. within a random process with a capital letter subscripted with its index, $X_{t}$. Realizations from these random variables will be denoted as $x_{t}$.
In our treatment of the linear MMSE solution, we assume wide-sense stationarity of the processes involved. A definition of wide-sense stationarity can be found in [26]. Stationarity is a necessary assumption to practically estimate the quantities needed to solve the Wiener-Hopf equations from realizations of a random process [19]. A strictly stationary random process is a random process where the joint distributions (not just the second-order moments) do not change with shifts in time, which is required in our nonlinear MMSE solution.
We now review the linear MMSE solution given in [21] (also see [27]). Let the space of square-integrable r.v.s. defined on $(\Omega,\mathcal{B},P)$ be denoted as $L^{2}(\Omega,\mathcal{A},P)$. This is the space of all r.v.s., $W$ such that $\|W\|_{2}=\int_{\Omega}|W|^{2}dP<\infty$. The linear span of a random process, $X$, in $L^{2}(\Omega,\mathcal{A},P)$ is the smallest subspace of $L^{2}(\Omega,\mathcal{A},P)$ containing $X$ [27]. We can define this set by first defining the linear manifold $L(X_{t},t\in T)$ as the set of all r.v.s. with the form $W=\sum_{i=1}^{n}a_{i}X_{t_{i}}$ with $a_{i}\in\mathbb{R}$, $t_{i}\in T$, and $n\in\mathbb{N}$. While $L(X_{t},t\in T)$ is a linear manifold, it is not complete. We can complete this space by including the limits of all Cauchy sequences of elements in $L(X_{t},t\in T)$. This complete set is the linear span of a random process in $L^{2}(\Omega,\mathcal{A},P)$, denoted as $L^{2}(X)$. Consider two r.v.s. $W,V\in L^{2}(X)$ where $W=\sum_{t\in T}a_{t}X_{t}$, $V=\sum_{s\in T}b_{s}X_{s}$, and $a_{t},b_{s}\in\mathbb{R}$. The inner product in $L^{2}(X)$ can be written as
$\langle W,V\rangle_{L^{2}}=\mathbb{E}[WV]=\mathbb{E}\left[\sum_{s,t\in T}a_{t}b_{s}X_{t}X_{s}\right]=\sum_{s,t\in T}a_{t}b_{s}\mathbb{E}[X_{t}X_{s}].$ (1)
Then with the positive semi-definite covariance function $R(s,t)=\mathbb{E}[X_{t}X_{s}]$, we see that the inner product between any two r.v.s. in $L^{2}(X)$ can be written in the RKHS whose kernel is defined by the covariance function of the random process $X$ as,
$\langle W,V\rangle_{L^{2}}=\sum_{s,t\in T}a_{t}b_{s}R(s,t)=\langle W^{{}^{\prime}},V^{{}^{\prime}}\rangle_{\mathcal{H}_{R}}.$ (2)
@@
Suppose we are given the random process $X$ as input and $Z$ as the desired random process. As a consequence of Hilbert projection theorem, the linear MMSE solution can be given as the projection of $Z$ into $H_{R}$. In [21], it is shown that the cross-covariance function is the linear MMSE solution in $\mathcal{H}_{R}$. Therefore, the MMSE solution is the inner product between the cross covariance function $\rho$ with $X$ in $\mathcal{H}_{R}$ that is,
$Z^{*}=\langle\rho,X\rangle_{\mathcal{H}_{R}}=\sum_{s,t\in T}X_{t}R(s,t)^{\dagger}\rho(s),\qquad\rho(s)=\mathbb{E}[ZX_{s}],$ (3)
where $\dagger$ is the Moore-Penrose pseudo inverse. In discrete time, this solution is equivalent to the Wiener solution, where the optimal impulse response is given by $w^{*}=\mathbf{R}^{\dagger}\bm{\rho}$, where $\mathbf{R}$ is the auto-covariance matrix and $\bm{\rho}$ is the cross-covariance vector. This demonstration shows that in the data-dependent RKHS, $\mathcal{H}_{R}$, there is no need to search for the linear MMSE solution. The solution is the cross-covariance function. The problem is that this solution is still linear in the input space. We now give the generalization of Parzen’s work in [21] to include nonlinear functions.